What are the three next terms in the sequence, 1, 1, 2, 3, 5
Found 2 solutions by htmentor, ikleyn:
Answer by htmentor(1343)

Looking over the pattern, it is clear that terms in the sequence are generated
by summing the two terms that precede them, i.e., 2=1+1, 3=2+1, and 5=3+2
Therefore, the next term will be 3+5=8, and so on.
The next 3 terms are 8, 13 and 21.

Answer by ikleyn(52777)

.
Fibonacci numbers.

https://en.wikipedia.org/wiki/Fibonacci_number

https://en.wikipedia.org/wiki/Fibonacci_number